A violet ray of light leaves a first medium at an angle of 0.79 radians to the normal and enters a second medium at 0 56 radians to the normal. Given that the refractive index of the second medium is 1.33. What is the refractive index of the first medium?

1
Expert's answer
2022-02-14T17:03:15-0500

"\frac{n_2}{n_1}=\frac{\sin\beta}{\sin\alpha},"

"n_2=n_1\frac{\sin\beta}{\sin\alpha}=1."